Using divisibility tests, determine which of the following numbers are divisible by 11: 
901153

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Sum of digits at odd places = 3 + 1 + 0 =4
Sum of digits at even places = 5 + 1 + 9 = 15
Difference = 15–4 = 11
Since, the difference between sum of digits at odd places and sum of digits at even places is 11 which is divisible by 11.
Hence, 901153 is divisible by 11

Thanks